OpenStudy (anonymous):

The graph of the following system of equations is 2x + y = 6 6x + 3y = 18 Overlapping lines Parallel lines Intersecting lines

sam (.sam.):

2x + y = 6 multiply 3 6x + 3y = 18 ---------------------------- 6x + 3y = 18 6x + 3y = 18 Both same equation so its overlapping lines